40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 304 of 344  Not logged in ELOG logo
ID Date Author Type Category Subject
  2024   Tue Sep 29 23:43:46 2009 robUpdateSUSITMY UL OSEM

We had a redo of elog entry 975 tonight.  The noisy OSEM was fixed by jiggling the rack end of the long cable.  Don't know exactly where--I also poked around the OSEM PD interface board.

In the attached PDF the reference trace is the noisy one.

  2023   Tue Sep 29 22:51:20 2009 KojiUpdateMZPossible gain mis-calibration at other places (Re: MZ work done)

Probably there is the same mistake for the PMC gain slider. Possibly on the FSS slider, too???

Quote:

2) The EPICS setting for the MZ gain slider was totaly wrong.
    Today I learned from the circuit, the full scale of the gain slider C1:PSL-MZ_GAIN gave us +/-10V at the DAC.
    This yield +/-1V to V_ctrl of the AD602 after the internal 1/10 attenuation stage.
    This +/-1V didn't correspond to -10dB~+30dB, but does -22dB~+42dB and is beyond the spec of the chip.

  2022   Tue Sep 29 21:51:32 2009 KojiUpdateMZMZ work done : some noise checking

The previous "+15" was Vctrl = 0.25 [V]. Which was +18 dB.

Quote:

Since we used to run with a gain slider setting of +15 dB on the MZ, I wanted to check that the new setting of +30dB was OK.

 

  2021   Tue Sep 29 21:37:09 2009 ranaUpdateMZMZ work done : some noise checking

Since we used to run with a gain slider setting of +15 dB on the MZ, I wanted to check that the new setting of +30dB was OK. It is.

To check it I turned it up and looked for some excess noise in the ISS or in the MC. There was none. I also set the input offset slider by unlocking the PMC and zeroing the mixer monitor readback. The new slider setting is -6.5V.

I don't know why we would need more gain on the MZ loop, but we can have some if we want it by turning up the gain before the servo (optical or RF). The attached plot shows the MC_F and ISS signals with the ISS loop on and off. There was no change in either of these spectra with the MZ gain high or low.

  2020   Tue Sep 29 18:21:41 2009 KojiUpdateMZMZ work done

The MZ work completed. I replaced the bad cross connection terminal. The gain slider is working now.

I looked at the error spectrum on an FFT analyzer. I could see the lock was more tight.

Then I proceeded to the MZ epics panel.

1) C1:PSL-MZ_MZTRANSPD has no meaning (not connected). So I put  C1:PSL-ISS_INMONPD as the MZ trans monitor.

2) The EPICS setting for the MZ gain slider was totaly wrong.
    Today I learned from the circuit, the full scale of the gain slider C1:PSL-MZ_GAIN gave us +/-10V at the DAC.
    This yield +/-1V to V_ctrl of the AD602 after the internal 1/10 attenuation stage.
    This +/-1V didn't correspond to -10dB~+30dB, but does -22dB~+42dB and is beyond the spec of the chip.

    The gain of AD602 is calculated by

G [dB] = 32 V_crtl + 10,  for -0.625 [V]< V_ctrl < +0.625 [V].

    In order to fix this I used the following commands which overrode the EPICS parameters.
    The tip of EGUF/EGUL is to know how much the gain (virtually) goes for the full scale of the DAC output. 

ezcawrite C1:PSL-MZ_GAIN.EGUF 42
ezcawrite C1:PSL-MZ_GAIN.EGUL -22
ezcawrite C1:PSL-MZ_GAIN.DRVH 30
ezcawrite C1:PSL-MZ_GAIN.DRVL -10
ezcawrite C1:PSL-MZ_GAIN.HOPR 30
ezcawrite C1:PSL-MZ_GAIN.LOPR -10

   and for the permanent change I modified the db file /cvs/cds/caltech/target/c1iool0/c1iooMZservo.db
   This will be active when cliool0 is rebooted.

# This yields the output limited to -6.25V ~ +6.25V, which corresponds to -10dB ~ +30dB
# modified by Koji Arai (29-Sept-2009)
grecord(ao,"C1:PSL-MZ_GAIN")
{
        field(DESC,"GAIN- overall pre-modecleaner servo loop gain")
        field(SCAN,"Passive")
        field(PINI,"YES")
        field(DISV,"1")
        field(DTYP,"VMIVME-4116")
        field(OUT,"#C3 S5 @")
        field(EGUF,"42")
        field(EGUL,"-22")
        field(PREC,"1")
        field(EGU,"dB")
        field(HOPR,"30")
        field(LOPR,"-10")
        field(DRVH,"30")
        field(DRVL,"-10")
        field(LINR,"LINEAR")
        field(OROC,"0")
        field(DOL,"0")
}

# previous code
grecord(ao,"C1:PSL-MZ_GAIN")
{
        field(DESC,"GAIN- overall pre-modecleaner servo loop gain")
        field(SCAN,"Passive")
        field(PINI,"YES")
        field(DISV,"1")
        field(DTYP,"VMIVME-4116")
        field(OUT,"#C3 S5 @")
        field(EGUF,"30")
        field(EGUL,"-10")
        field(PREC,"4")
        field(EGU,"Volts")
        field(HOPR,"30")
        field(LOPR,"-10")
        field(LINR,"LINEAR")
        field(OROC,"0")
        field(DOL,"0")
}

Quote:

12:45 I started the work on MZ. Thus the MZ was unlocked.

Fond the bad connection on the FLKM 64pin cross connection board. We need the replacement.

I went to Wilson and got the replacement, two VME extender boards, three 7815, and three 7915. Thanks, Ben!

 

  2019   Tue Sep 29 16:14:44 2009 robConfigurationElectronicsRob is breaking stuff....

Quote:

Koji and I were looking for an extender card to aid with MZ board testing.  Rob went off on a quest to find one.  He found 2 (in addition to the one in the drawer near the electronics bench which says "15V shorted"), and put them in some empty slots in 1X1 to test them out.  Somehow, this burned a few pins on each board (1 pin on one of them, and 3 pins on the other). We now have 0 functioning extender cards: unfortunately, both extender cards now need fixing.  The 2 slots that were used in 1X1 now have yellow electrical tape covering the connectors so that they do not get used, because the ends of the burnt-off pins may still be in there. 

In other, not-Rob's-fault news, the Martian network is down...we're going to try to reset it so that we have use of the laptops again.

 

This happened when I plugged the cards into a crate with computers, which apparently is a no-no.  The extender cards only go in VME crates full of in-house, LIGO-designed electronics.

  2018   Tue Sep 29 12:47:08 2009 KojiUpdateMZMZ unlocked

12:45 I started the work on MZ. Thus the MZ was unlocked.

Found the bad connection on the FLKM 64pin cross connection board. We need a replacement.

I went to Wilson and got the replacement, two VME extender boards, three 7815, and three 7915. Thanks, Ben!

  2017   Tue Sep 29 10:44:29 2009 KojiUpdateMZMZ investigation

Rana, Jenne, Koji

Last night we checked MZ. The apparent thing we found was the gain slider does not work.
The slider actually changes the voltage at the cross connection of 1Y2 (31 pin4?), the gain does not change.
The error spectrum didn't change at all even when the slider was moved.

Rana poked the flat cable at the bottom of 1Y2, we had no imporvement.

We coudn't find the VME extender board, so we just replaced AD602 (=VGA) and LT1125 (=Buffer for the ctrl voltage).
Even after the replacement, the gain slider is not working yet.

Today, I will put a lead or probe to the board to see whether the slider changes the voltage on the board or not.

Somehow the gain is sitting at a intermediate place that is not to low not to high. So I still don't know the gain slider
is the cause of the MZ instability or not.

  2016   Tue Sep 29 01:50:10 2009 robConfigurationLSCnew modulation frequencies

Mode cleaner length measured tonight.

 

33196198

132784792

165980990

199177188

[Tag by KA: modulation frequency, MC length]

  2015   Mon Sep 28 23:44:18 2009 KojiOmnistructureSAFETYCrappy power outlet

Jenne, Koji

Tonight we found that the wireless for Martian network was down.
We inspected the router and found the power was down. The power of the weather station was also down.

By touching the power outlet which they are connected, the power changes on and off.
This problematic power outlet has a label "L#17" just below the photograph of the mk I (1989).
The plug was connected to the left one.

As it was scary, we moved the power plug to the next one (L#19).
The wireless router and the weather station were powered now,
though the weather station is showing a wrong time in its clock.

  2014   Mon Sep 28 23:13:14 2009 JenneConfigurationElectronicsRob is breaking stuff....

Koji and I were looking for an extender card to aid with MZ board testing.  Rob went off on a quest to find one.  He found 2 (in addition to the one in the drawer near the electronics bench which says "15V shorted"), and put them in some empty slots in 1X1 to test them out.  Somehow, this burned a few pins on each board (1 pin on one of them, and 3 pins on the other). We now have 0 functioning extender cards: unfortunately, both extender cards now need fixing.  The 2 slots that were used in 1X1 now have yellow electrical tape covering the connectors so that they do not get used, because the ends of the burnt-off pins may still be in there. 

In other, not-Rob's-fault news, the Martian network is down...we're going to try to reset it so that we have use of the laptops again.

  2013   Mon Sep 28 17:39:34 2009 robUpdatePSLproblems

The PSL/IOO combo has not been behaving responsibly recently. 

The first attachment is a 15 day trend of the MZ REFL, ISS INMON, and MC REFL power.  These show two separate problems--recurring MZ flakiness, which may actually be a loose cable somewhere which makes the servo disengage.  Such disengagement is not as obvious with the MZ as it is with other systems, because the MZ is relatively stable on its own.  The second problem is more recent, just starting in the last few days.  The MC is drifting off the fringe, either in alignment, length, or both.  This is unacceptable.

The second attachment is a two-day trend of the MC REFL power.  Last night I carefully put the beam on the center of the MC-WFS quads.  This appears to have lessened the problem, but it has not eliminated it. 

It's probably worth trying to re-measure the MCWFS system to make sure the control matrix is not degenerate. 

  2012   Mon Sep 28 11:52:23 2009 JenneUpdateTreasureOAF screen added to the screenshots webpage

I used Kakeru's instructions in elog 1221 to add the C1OAF screen (still called C1ASS_TOP) to the medm screenshots webpage.  The tricky part of this is figuring out that the file that needs editing is in fact in /cvs/cds/projects/statScreen, not /cvs/cds/caltech/statScreen, as claimed in the entry. 

  2011   Mon Sep 28 02:24:05 2009 ranaUpdateLockingMC1/3 Dewhitening found OFF: Turned back ON

While trying to make the OAF work, I found that the XYCOM switches for MC1/3 have been set in the bad way for awhile. This means that the hardware filters were bypassed and that MC1 & MC3 were moving around too much at high frequency and possibly causing trouble with the locking. I have put them back into the default position.

On Friday, Jenne and I were playing around with turning the dewhitening off/on to see if it efffected the OAF stability. At the time, I didn't pay too much attention to what the state was. Looks like it was in the wrong state (hardware bypassed) when we found it. For the OAF work, we generally want it in that bypassed state, but its bad because it makes noise in the interferometer. The bits in question are bits 16-23 on the XYCOM screen.

I have updated the snapshot and set the screen in the appropriate settings. I used a swept sine measurement to verify the filter state. In the attached plot, green corresponds to XYCOM green and red corresponds to red.

  2010   Sun Sep 27 23:21:14 2009 ranaUpdatePSLSLOWscan result

Quote:

What I like to try:
a) Change the NPRO temp to more cold side.
b) Change the MOPA head temp to a bit hot side.
c) Tweak the MOPA current (is it difficult?)

 I think that the AMPMON ND problem was just that the responsivity changes with angle. So when I aligned it a little we got some few% improvement in the signal which is not a real power increase.

I don't think we can adjust any of the MOPA parameters because the controller is broken, but we can try the NPRO crystal temperature.

  2009   Sun Sep 27 15:25:58 2009 KojiUpdatePSLSLOWscan result
Oh, AMPMON dependence could be an artifact of the ND filter???
For my case, it should be real dependence on the NPRO wavelength,
as the other PDs like the PMC reflection (PMC_RFPDDC) and the RC reflection (FSS_RFPDDC) show the same dependence.
  2008   Sun Sep 27 14:45:45 2009 KojiUpdatePSLSLOWscan result

I ran (script dir)/PSL/FSS/SLOWscan on op440m from 11:30 to 12:30 on 27th. Although Rana and later I myself set "timed bombs" for the scan, they did not work as they have probably been ran on Linux. After the scan I relocked PMC, FSS, and MZ . MC locked automatically.

Observation:

1. To keep away from the mode hop, FSS_SLOWDC is to be at around 0. The values -5 ~ -6 is the place for the power, which is my preference for now. BTW, the mode hop only appears to the PSL output (=AMPMON) is this normal?

2. The PSL output looks dependent on the NPRO wavelength. The NPRO output and the PSL output tends to be high when the FSS_SLOWDC is low (= LTMP: Laser Crystal Temp is low). Also there is a step at the LTMP where we think the mode hop is present. This may cause the daily PSL output variation which induced by the daily change of the reference cavity length.

My naive speculation is that the NPRO wavelength is too long (= hot side) for the MOPA absorption as the MOPA heads are cooled to 19deg.

3. Scanning of -10 to +10 changes the LTMP from 42-49deg. This is almost 1/10 of the NPRO capability. The manual told us that we should be able to scan the crystal temperature +/-16deg (about 30deg to 60deg).

What I like to try:
a) Change the NPRO temp to more cold side.
b) Change the MOPA head temp to a bit hot side.
c) Tweak the MOPA current (is it difficult?)

  2007   Sun Sep 27 12:52:56 2009 ranaUpdateMOPAIncreasing the power from the MOPA

This is a trend of the last 20 days. After our work with the NPRO, we have recovered only 5% in PMC trans power, although there's an apparent 15% increase in AMPMON.

The AMPMON increase is partly fake; the AMPMON PD has too much of an ND filter in front of it and it has a strong angle dependence. In the future, we should not use this filter in a permanent setup. This is not a humidity dependence.

The recovery of the refcav power mainly came from tweaking the two steering mirrors just before and just after the 21.5 MHz PC. I used those knobs because that is the part of the refcav path closest to the initial disturbance (NPRO).

BTW, the cost of a 1W Innolight NPRO is $35k and a 2W Innolight NPRO is $53k. Since Jenne is on fellowship this year, we can afford the 2W laser, but she has to be given priority in naming the laser.

  2006   Sat Sep 26 13:55:20 2009 JenneUpdateMZMZ was locked in a bad place

I found the MZ locked in a bad place earlier today.  It was locked in a similarly bad spot yesterday after we fixed the cable situation for 126MOPA_126MON, with reflection of ~0.8, rather than the nominal 0.305.  It's good now though. 

  2005   Fri Sep 25 19:56:08 2009 ranaConfigurationComputersNTPD restarted on c1dcuepics (to fix the MEDM screen times)
restarted ntp on op440m using this syntax
>su
>/etc/init.d/xntpd start -c /etc/inet/ntp.conf

gettting the time on scipe25 (for the MEDM screen time) working was tougher. The /etc/ntp.conf file was pointing
to the wrong server. Our NAT / Firewall settings require some of our internal machines to go through the gateway
to get NTPD to work. Curiously, some of the linux workstations don't have this issue.
The internal network machines should all have the same file as scipe25's /etc/ntp.conf:

server nodus

and here's how to check that its working:

[root@c1dcuepics sbin]# ./ntpq -p
     remote           refid      st t when poll reach   delay   offset  jitter
==============================================================================
 nodus.ligo.calt 0.0.0.0         16 u    -   64    0    0.000    0.000 4000.00
*nodus.ligo.calt usno.pa-x.dec.c  2 u   29   64  377    1.688  -65.616   6.647
-lime7.adamantsy clock.trit.net   3 u   32   64  377   37.448  -72.104   4.641
-montpelier.ilan .USNO.           1 u   19   64  377   18.122  -74.984   8.305
+spamd-0.gac.edu nss.nts.umn.edu  3 u   28   64  377   72.086  -66.787   0.540
-mighty.poclabs. time.nist.gov    2 u   30   64  377   71.202  -61.127   4.067
+monitor.xenscal clock.sjc.he.ne  2 u   16   64  377   11.855  -67.105   6.368
  2004   Fri Sep 25 19:55:59 2009 JenneUpdateAdaptive FilteringSubtraction of the microseism using Adaptive Filtering!

[Rana, Jenne]

The OAF system did something useful today!  Attached is a plot.  Black is the reference (13 averages) with the OAF off.  Blue is the output of the OAF, and red is the reduced MC_L signal (13 averages).  If you turn tau and mu both to 0, it "pauses" the filter, but keeps the feedforward system working, so that you can take a long average to get a better idea of how well things are working. If you ramp down the output of the CORR filter bank, that lets you take a long average with the OAF "off", but doesn't mess up your nicely adapted filter.  The cyan and gold traces in the upper plot are 2 of the Guralp channels, so you can see the real seismic motion.

In the lower plot, you can see that the cyan and light green seismic channels have good coherence with IOO-MC_L (the names don't really mean anything right now...these 2 seismometer channels are the 2 Guralps' channels, one per end of the MC, which are aligned with the MC.)  The dark blue trace is the coherence between IOO-MC_L and the output of the OAF.

500 taps, delay=5, 2 Guralp channels (the ones aligned with the MC), tau~0.00001 (probably), and mu~0.01 or 0.005

  2003   Fri Sep 25 17:51:51 2009 KojiUpdateMOPASolved (Re: Total MOPA power is constant, but the NPRO's power has decreased after last night's activities?)

Jenne, Koji

The cause of the decrease was found and the problem was solved. We found this entry, which says

Yoich> We opened the MOPA box and installed a mirror to direct a picked off NPRO beam to the outside of the box through an unused hole.
Yoich> We set up a lens and a PD outside of the MOPA box to receive this beam. The output from the PD is connected to the 126MON cable.

We went to the PSL table and found the dc power cable for 126MOPA_AMPMON was clipping the 126MON beam.
We also made a cable stay with a pole and a cable tie.

After the work, 126MON went up to 161 which was the value we saw last night.


We also found that the cause of the AMPMON signal change by the DAQ connection, mentioned in this entry:

Jenne> 6.  We teed off of the AMPMON photodiode so that we could see the DC values on a DMM. 
Jenne> When we used a T to connect both the DMM and the regular DAQ cable, the DMM read
Jenne> a value a factor of 2 smaller than when the DMM was connected directly to the PD.

We found a 30dB attenuator is connected after the PD. It explains missing factor of 2.

Quote:

[Koji, Jenne]

Steve pointed this out to me today, and Koji and I just took a look at it together:  The total power coming out of the MOPA box is constant, about 2.7W.  However, the NPRO power (as measured by 126MOPA_126MON) has decreased from where we left it last night.  It's an exponential decay, and Koji and I aren't sure what is causing it.  This may be some misalignment on the PD which actually measures 126MON or something though, because 126MOPA_LMON, which measures the NPRO power inside the NPRO box (that's how it looks on the MEDM screen at least...) has stayed constant.  I'm hesitant to be sure that it's a misalignment issue since the decay is gradual, rather than a jump. 

Koji and I are going to keep an eye on the 126MON value.  Perhaps on Monday we'll take a look at maybe aligning the beam onto this PD, and look at the impedance of both this PD, and the AMPMON PD to see why the reading on the DMM changed last night when we had the DAQ cable T-ed in, and not T-ed in. 

 

  2002   Fri Sep 25 16:45:29 2009 JenneUpdateMOPATotal MOPA power is constant, but the NPRO's power has decreased after last night's activities?

[Koji, Jenne]

Steve pointed this out to me today, and Koji and I just took a look at it together:  The total power coming out of the MOPA box is constant, about 2.7W.  However, the NPRO power (as measured by 126MOPA_126MON) has decreased from where we left it last night.  It's an exponential decay, and Koji and I aren't sure what is causing it.  This may be some misalignment on the PD which actually measures 126MON or something though, because 126MOPA_LMON, which measures the NPRO power inside the NPRO box (that's how it looks on the MEDM screen at least...) has stayed constant.  I'm hesitant to be sure that it's a misalignment issue since the decay is gradual, rather than a jump. 

Koji and I are going to keep an eye on the 126MON value.  Perhaps on Monday we'll take a look at maybe aligning the beam onto this PD, and look at the impedance of both this PD, and the AMPMON PD to see why the reading on the DMM changed last night when we had the DAQ cable T-ed in, and not T-ed in. 

  2001   Fri Sep 25 16:10:17 2009 JenneUpdateAdaptive FilteringSome progress on OAF, but more still to be done

[Jenne, Sanjit]

It seems now that we are able to get the OAF system to do a pretty good job of approximating the MC_L signal, but we can't get it to actually do any subtracting.  I think that we're not correctly setting the phase delay between the witness and the MC_L channels or something (I'm not sure though why we get a good filter match if the delay is set incorrectly, but we do get a good filter match for very different delay settings: 1, 5, 100, 1000 all seem to do equally well at adjusting the filter to match MC_L). 

The Matt Evans document in elog 395 suggests measuring the phase at the Nyquist frequency, and calculating the appropriate delay from that.  The sticking point with this is that we can't get test points for any channel which starts with C1:ASS.  I've emailed Alex to see what he can do about this.  Elog 1982 has a few words about how we're perhaps using a different awgtpman on the ass machine than we used to, which may be part of the problem. 

The golden plan, which in my head will work perfectly, is as follows: Alex will fix the testpoint problem, then Sanjit and I will be able to measure the phase between our OAF signal and the incoming MC_L signal, we will be able to match them as prescribed in the Matt Evans document, and then suddenly the Adaptive Filtering system will do some actual subtracting!

The plot below shows the Reference MC_L without any OAF system (black), the output of the OAF (green), and the 'reduced' MC_L (red).  As you can see, the green trace is doing a pretty good job of matching the black one, but the red trace isn't getting reduced at all.

  2000   Thu Sep 24 21:04:15 2009 JenneUpdateMOPAIncreasing the power from the MOPA

[Jenne, Rana, Koji]

Since the MOPA has been having a bad few weeks (and got even more significantly worse in the last day or so), we opened up the MOPA box to increase the power.  This involved some adjusting of the NPRO, and some adjusting of the alignment between the NPRO and the Amplifier.  Afterward, the power out of the MOPA box was increased.  Hooray! 

Steps taken:

0.  Before we touched anything, the AMPMON was 2.26, PMC_Trans was 2.23, PSL-126MOPA_126MON was 152 (and when the photodiode was blocked, it's dark reading was 23).

1.  We took off the side panel of the MOPA box nearest the NPRO, to gain access to the potentiometers that control the NPRO settings.  We selectively changed some of the pots while watching PSL-126MOPA_126MON on Striptool.

2.  We adjusted the pot labeled "DTEMP" first. (You have to use a dental mirror to see the labels on the PCB, but they're there). We went 3.25 turns clockwise, and got the 126MON to 158. 

3. To give us some elbow room, we changed the PSL-126MOPA_126CURADJ from +10.000 to 0.000 so that we have some space to move around on the slider.  This changed 126MON to 142. The 126MOPA_CURMON was at 2.308.

4.  We tried adjusting the "USR_CUR" pot, which is labeled "POWER" on the back panel of the NPRO (you reach this pot through a hole in the back of the NPRO, not through the side which we took off, like all the other pots today).  This pot did nothing at all, so we left it in its original position.  This may have been disabled since we use the slider.

5.  We adjusted the CUR_SET pot, and got the 126MON up to 185.  This changed the 126MOPA_CURMON to 2. 772 and the AMPMON to 2.45

We decided that that was enough fiddling with the NPRO, and moved on to adjusting the alignment into the Amplifier.

6.  We teed off of the AMPMON photodiode so that we could see the DC values on a DMM.  When we used a T to connect both the DMM and the regular DAQ cable, the DMM read a value a factor of 2 smaller than when the DMM was connected directly to the PD.  This shouldn't happen.....it's something on the to-fix-someday list.

7.  Rana adjusted the 2 steering mirrors immediately in front of the amplifier, inside the MOPA box.  This changed the DMM reading from its original 0.204 to 0.210, and the AMPMON reading from 2.45 to 2.55. While this did help increase the power, the mirrors weren't really moved very much.

8.  We then noticed that the beam wasn't really well aligned onto the AMPMON PD.  When Rana leaned on the MOPA box, the PD's reading changed.  So we moved the PD a little bit to maximize its readings.  After this, the AMPMON read 2.68, and the DMM read 0.220.

9.  Then Rana adjusted the 2 waveplates in the path from the NPRO to the Amplifier.  The first waveplate in the path didn't really change anything.  Adjusting the 2nd waveplate gave us an AMPMON of 2.72, and a DMM reading of 0.222.

10.  We closed up the MOPA box, and locked the PMC.  Unfortunately, the PMC_Trans was only 1.78, down from the 2.26 when we began our activities.  Not so great, considering that in the end, the MOPA power went up from 2.26 to 2.72.

11.  Koji and I adjusted the steering mirrors in front of the PMC, but we could not get a transmission higher than 1.78.

12.  We came back to the control room, and changed the 126MOPA_126CURADJ slider to -2.263 which gives a 126MOPA_CURMON to 2.503.  This increased PMC_TRANS up to 2.1. 

13.  Koji did a bit more steering mirror adjustment, but didn't get any more improvement.

14.  Koji then did a scan of the FSS SLOW actuator, and found a better temperature place (~ -5.0)for the laser to sit in.  This place (presumably with less mode hopping) lets the PMC_TRANS get up to 2.3, almost 2.4.  We leave things at this place, with the 126MOPA_126CURADJ slider at -2.263. 

Now that the MOPA is putting out more power, we can adjust the waveplate before the PBS to determine how much power we dump, so that we have ~constant power all the time.

 

Also, the PMCR view on the Quad TVs in the Control Room has been changed so it actually is PMCR, not PMCT like it has been for a long time.

  1999   Thu Sep 24 20:17:05 2009 ranaSummaryLSCComparison of Material Properties for the new RFPD Mounts
  Steel Brass Aluminum Delrin
Density (kg/m^3) 7850 8500 2700 1420
CTE (ppm/C) 12 19 23 100

Young's

Modulus

(GPa)

200 110 69 2
Hardness        
Color grey gold light silver any

 

  1998   Thu Sep 24 19:35:20 2009 ranaHowToPhotos40m Google account

I've created a 40m Google account. Please post all the 40m related photos to this site. If you don't already have it, download Picasa to make this easier.

40m Installation Photos">

the password is in the usual password place.

  1997   Thu Sep 24 15:45:27 2009 robUpdateIOOMC OLG

I measured the mode cleaner open loop gain with the HP3563A

The UGF is 64kHz, phase margin is 28 deg.

  1996   Wed Sep 23 20:02:11 2009 JenneAoGComputersGremlins in the RFM

Quote:

A cosmic ray struck the RFM in the framebuilder this afternoon, causing hours of consternation.  The whole FE system is just now coming back up, and it appears the mode cleaner is not coming back to the same place (alignment).

 

rob, jenne

 Jenne, Rana, Koji

The mode cleaner has been realigned, using a combination of techniques.  First, we used ezcaservo to look at C1:SUS-MC(1,3)_SUS(DOF)_INMON and drive C1:SUS-MC(1,3)_(DOF)_COMM, to put the MC1 and MC3 mirrors back to their DriftMon values.  Then we looked at the MC_TRANS_SUM on dataviewer and adjusted the MC alignment sliders by hand to maximize the transmission.  Once the transmission was reasonably good, we saw that the spot was still a little high, and the WFS QPDs weren't centered.  So Koji and I went out and centered the WFS, and now the MC is back to where it used to be.  The MC_TRANS QPD looks nice and centered, so the pointing is back to where it used to be.

  1995   Wed Sep 23 19:36:41 2009 ranaUpdatePSLRC temperature performance

This first plot shows the RC temperature channels' performance from 40 days ago, before we disabled the MINCO PID controller. Although RCTEMP is supposed to be the out of loop sensor, what we really care about is the cavity length and so I've plotted the SLOW. To get the SLOW on the same scale, I've multiplied the channel by 10 and then adjusted the offset to get it on the same scale.

 The second plot shows a period after that where there is no temperature control of the can at all. Same gain scaling has been applied to SLOW as above, so that instead of the usual 1 GHz/V this plot shows it in 0.1 GHz/V.

The third plot shows it after the new PID was setup.

Summary: Even though the PID loop has more gain, the true limit to the daily fluctuations in the cavity temperature and the laser frequency are due to the in-loop sensors measuring the wrong thing. i.e. the out-of-loop temperature is too different from the in-loop sensor. This can possibly be cured with better foam and better placement of the temperature sensors. Its possible that we're now just limited by the temperature gradients on the can.

  1994   Wed Sep 23 17:32:37 2009 robAoGComputersGremlins in the RFM

A cosmic ray struck the RFM in the framebuilder this afternoon, causing hours of consternation.  The whole FE system is just now coming back up, and it appears the mode cleaner is not coming back to the same place (alignment).

 

rob, jenne

  1993   Fri Sep 18 16:26:02 2009 steveSummaryPSLNeslab chiller is OK

Rob found puddles of water very close to the chiller during lunch time. We raised the unit and took the side cover off. All surfaces were dry and the water level in the tub normal.

Later on we discovered that one of the Vons distilled water bottle was leaking. Jenne and I checked for excess amount of condensing water droplets inside the MOPA box.

On the bare,not insulated tubing and valve are loaded with droplets of water. Relative humidity is 44% at 24 C and HEPA filter speed set to 80 V in the enclosure.

 

  1992   Fri Sep 18 16:05:08 2009 JenneOmnistructurePSLwater under the laser chiller

Quote:

rob, koji, steve

We noticed some water (about a cup) on the floor under the NESLAB chiller today.  We put the chiller up on blocks and took off the side panel for a cursory inspection, but found no obvious leaks.  We'll keep an eye on it.

 The culprit has been found:  One of the bottles of chiller water had a tiny leak in it, and apparently the floor is sloped just right to make it look like the water had been coming from under the chiller.  All is well again in the world of chilled water.

  1991   Fri Sep 18 14:25:00 2009 robOmnistructurePSLwater under the laser chiller

rob, koji, steve

We noticed some water (about a cup) on the floor under the NESLAB chiller today.  We put the chiller up on blocks and took off the side panel for a cursory inspection, but found no obvious leaks.  We'll keep an eye on it.

  1990   Thu Sep 17 15:05:47 2009 robUpdateComputersawgtpman on c1omc failing to start

Quote:

[root@c1omc controls]# /opt/gds/awgtpman -2 &
[1] 16618
[root@c1omc controls]# mmapped address is 0x55577000
32 kHz system
Spawn testpoint manager
no test point service registered
Test point manager startup failed; -1

[1]+  Exit 1                  /opt/gds/awgtpman -2

 

 

 

 

 

 

This turned out to be fallout from the /cvs/cds transition.  Remounting and restarting fixed it.

  1989   Thu Sep 17 14:17:04 2009 robUpdateComputersawgtpman on c1omc failing to start

[root@c1omc controls]# /opt/gds/awgtpman -2 &
[1] 16618
[root@c1omc controls]# mmapped address is 0x55577000
32 kHz system
Spawn testpoint manager
no test point service registered
Test point manager startup failed; -1

[1]+  Exit 1                  /opt/gds/awgtpman -2

 

 

 

  1988   Wed Sep 16 11:58:11 2009 JenneUpdateAdaptive FilteringNew Filters for Adaptive Filtering

When Sanjit and I were looking at the adaptive filtering system on Monday and Friday, we noticed that turning on the Accelerometers (which had been used in the past) seemed to do good things, but that turning on the seismometers (which I just put into the system last week) made the OAF output integrate up.  Rana pointed out that this is an indication of a missing high pass filter.  And indeed, when I put the seismometers in, I neglected to copy the high pass filter at low frequencies, and the low pass at 64Hz from the accelerometer path to the seismometer path.  The accelerometers had a HP at 1Hz, which is okay since they don't really do useful things down to the mHz level.  I gave all of the seismometers HP at 1mHz.  These are now in the filter banks in the ASS_TOP_PEM screen.  The accelerometers are on channels 15, 16, 17, 18, 19, 20 and the seismometers are on channels 2, 3, 4, 10, 11, 12, 24.

I now need to modify the upass script to turn these filters on before doing adaptive filtering.

  1987   Tue Sep 15 15:46:05 2009 steveUpdatePEMPEM and VAC

FSS_RMTEMP is moving up and  daily fluctuations  are  less . 120 and 16 days plots are below.

  1986   Sat Sep 12 15:40:15 2009 ranaUpdatePSLRC response v. can temperature

I stepped the RC can temperature to see the response in the laser frequency. This gives a true measure of the thermal time constant of the RC. Its ~4 hours.

Since the RCPID screen now has a setpoint field, I can remotely type in 1 deg steps. The NPRO SLOW actuator locks the NPRO to the RC at long time scales and so we can use C1:PSL-FSS_SLOWDC to measure the RC length. By knowing what the step response time constant is, we can estimate the transfer function from can temperature to frequency noise and thereby make a better heater circuit.

 Does the observed temperature shift make any sense? Well, John Miller and I measured the SLOW calibration to be 1054 +/- 30 MHz / V.

We know that the thermal expansion coefficient of fused silica, alpha = 5.5 x 10-7 (dL/L)/deg. So the frequency shift ought to be alpha * c / lambda = 155 MHz / deg.

Instead we see something like 110 MHz / deg. We have to take more data to see if the frequency shift will actually asymptote to the right value or not. If it doesn't, one possibility is that we are seeing the effect of temperature on the reflection phase of the mirror coatings through the dn/dT and the thermal expansion of the dielectric layers. I don't know what these parameters are for the Ta2O5 layers.

A more useful measure of the frequency noise can be gotten by just looking at the derivative in the first 30 minutes of the step, since that short time scale is much more relevant for us. Its 0.04 V / hour / (2 deg) =>  860 (Hz/s)/deg.

In the frequency domain this comes out to be dnu/dT = 860 Hz/deg @ 0.16 Hz or dnu/dT = 137 *(1/f) Hz / deg.

Our goal for the reference cavity frequency noise is 0.01 * (1/f) Hz/rHz. So the temperature noise of the can needs to be < 0.1 mdeg / rHz.

  1985   Fri Sep 11 17:11:15 2009 SanjitUpdateASSOAF: progress made

[Jenne & Sanjit]

Good news: We could successfully send filtered output to MC1 @ SUS.

We used 7 channels (different combinations of 3 seismometer and six accelerometer)

We tried some values of \mu (0.001-0.005) & gain on SUS_MC1_POSITION:MCL and C1ASS_TOP_SUS_MC1 (0.1-1).

C1:ASS-TOP_SUS_MC1_INMON is huge (soon goes up to few times 10000), so ~0.1 gains at two places bring it down to a reasonable value.

Bad news: no difference between reference and filtered IOO-MC_L power spectra so far.

Plan of action: figure out the right values of the parameters (\mu, \tau, different gains, and may be some delays), to make some improvement to the spectra.

 ** Rana: there's no reason to adjust any of the MCL gains. We are not supposed to be a part of the adaptive algorithm.

  1984   Fri Sep 11 17:07:45 2009 JenneUpdateAdaptive FilteringMinor changes to ASS_TOP_PEM screen.

There was some uncertainty as to which channels were being input into the Adaptive Filtering screen, so I checked it out to confirm.  As expected, the rows on the ASS_TOP_PEM screen directly correspond to the BNC inputs on the PEM_ADCU board in the 1Y6 (I think it's 6...) rack.  So C1:ASS-TOP_PEM_1_INMON corresponds to the first BNC (#1) on the ADCU, etc. 

After checking this out, I put text tags next to all the inputs on the ASS_TOP_PEM screen for all of the seismometers (which had not been there previously).  Now it's nice and easy to select which witness channels you want to use for the adaptation.

  1983   Thu Sep 10 18:25:15 2009 ranaUpdatePSLc1psl rebooted for new RCPID database settings

I added a new database record (C1:PSL-FSS_RCPID_SETPOINT) to allow for changing of the RC setpoint while the loop is on. This will enable us to step the can's temperature and see the result in the NRPO's SLOWDC.

 

  1982   Thu Sep 10 17:47:25 2009 JenneUpdateComputerschanges to the startass scripts

[Rana, Jenne]

While I was mostly able to restart the c1ass computer earlier today, the filter banks were acting totally weird.  They were showing input excitations when we weren't putting any, and they were showing that the outputs were all zero, even though the inputs were non-zero and the input and the output were both enabled. The solution to this ended up being to use the 2nd to last assfe.rtl backup file.  Rana made a symbolic link from assfe.rtl to the 2nd to last backup, so that the startup.cmd script does not need to be changed whenever we alter the front end code.

The startup_ass script, in /caltech/target/gds/ which, among other things, starts the awgtpman was changed to match the instructions on the wiki Computer Restart page.  We now start up the /opt/gds/awgtpman .  This may or may not be a good idea though, since we are currently not able to get channels on DTT and Dataviewer for the C1:ASS-TOP_PEM channels.  When we try to run the awgtpman that the script used to try to start ( /caltech/target/gds/bin/ ) we get a "Floating Exception". We should figure this out though, because the /opt/gds/awgtpman does not let us choose 2kHz as an option, which is the rate that the ASS_TOP stuff seems to run at.

The last fix made was to the screen snapshot buttons on the C1:ASS_TOP screen.  When the screen was made, the buttons were copied from one of the other ASS screens, so the snapshots saved on the ASS_TOP screen were of the ASS_PIT screen.  Not so helpful.  Now the update snapshot button will actually update the ASS_TOP snapshot, and we can view past ASS_TOP shots.

  1981   Thu Sep 10 15:55:44 2009 JenneUpdateComputersc1ass rebooted

c1ass had not been rebooted since before the filesystem change, so when I was sshed into c1ass I got an error saying that the NFS was stale.  Sanjit and I went out into the cleanroom and powercycled the computer.  It came back just fine.  We followed the instructions on the wiki, restarting the front end code, the tpman, and did a burt restore of c1assepics. 

  1980   Wed Sep 9 10:13:31 2009 steveBureaucracyGeneralthe use of FAX machine, scanner & toaster

 

 You said that the use of FAXST was forbidden for phds and graduate students. I had to swear on the promise of not ever buying an other FAXST

  1979   Tue Sep 8 20:25:03 2009 JenneOmnistructureDMFDMF restarted

Quote:

I (think I) restarted DMF.  It's on Mafalda, running in matlab (not the complied version which Rana was having trouble with back in the day).  To start Matlab, I did "nohup matlab", ran mdv_config, then started seisBLRMS.m running.  Since I used nohup, I then closed the terminal window, and am crossing my fingers in hopes that it continues to work.  I would have used Screen, but that doesn't seem to work on Mafalda.

 Just kidding. That plan didn't work.  The new plan: I started a terminal window on Op540, which is ssh-ed into Mafalda, and started up matlab to run seisBLRMS.  That window is still open. 

Because Unix was being finicky, I had to open an xterm window (xterm -bg green -fg black), and then ssh to mafalda and run matlab there.  The symptoms which led to this were that even though in a regular terminal window on Op540, ssh-ed to mafalda, I could access tconvert, I could not make gps.m work in matlab.  When Rana ssh-ed from Allegra to Op540 to Mafalda and ran matlab, he could get gps.m to work.  So it seems like it was  a Unix terminal crazy thing. Anyhow, starting an xterm window on Op540m and ssh-ing to mafalda from there seemed to work.

Hopefully this having a terminal window open and running DMF will be a temporary solution, and we can get the compiled version to work again soon.

  1978   Tue Sep 8 20:15:33 2009 rana, jenneSummaryPSLRC temperature servo: Heater Voltage noise

We measured the voltage noise of the heater used to control the RC can temperature. It is large.

TEK00074.PNG

The above scope trace shows the voltage directly on the monitor outputs of the heater power supply. The steps are from the voltage resolution of the 4116 DAC.

We also measured the voltage noise on the monitor plugs on the front panel. If these are a true representation of the voltage noise which supplies the heater jacket, then we can use it to estimate the temperature fluctuations of the can. Using the spectrum of temperature fluctuations, we can estimate the actual length changes of the reference cavity.

I used the new fax/scanner/toaster that Steve and Bob both love to scan this HP spectrum analyzer image directly to a USB stick! It can automatically make PDF from a piece of paper.

The pink trace is the analyzer noise with a 50 Ohm term. The blue trace is the heater supply with the servo turned off. With the servo on (as in the scope trace above) the noise is much much larger because of the DAC steps.

  1977   Tue Sep 8 19:36:52 2009 JenneOmnistructureDMFDMF restarted

I (think I) restarted DMF.  It's on Mafalda, running in matlab (not the complied version which Rana was having trouble with back in the day).  To start Matlab, I did "nohup matlab", ran mdv_config, then started seisBLRMS.m running.  Since I used nohup, I then closed the terminal window, and am crossing my fingers in hopes that it continues to work.  I would have used Screen, but that doesn't seem to work on Mafalda.

  1976   Tue Sep 8 19:30:33 2009 ranaUpdatePSLc1psl rebooted for new RCPID database settings

The RC thermal PID is now controllable from its own MEDM screen which is reachable from the FSS screen. The slowpid.db and psl.db have been modified to add these records and all seems to be working fine.

Also, I've attached the c1psl startup output that we got on the terminal. This is just for posterity.

I'm also done tuning the PID for now. Using Kp = -1.0, Ki = -0.01, and Kd = 0, the can servo now has a time constant of ~10 minutes and good damping as can be seen in the StripTool snap below. These values are also now in the saverestore.req so hopefully its fully commissioned.

I bet that its much better now than the MINCO at holding against the 24 hour cycle and can nicely handle impulses (like when Steve scans the table). Lets revisit this in a week to see if it requires more tuning.

  1975   Tue Sep 8 17:57:30 2009 JenneUpdatePEMAll the Acc/Seis working again

All of the accelerometers and seismometers are plugged in and functional again.  The cables to the back of the accelerometer preamp board (sitting under the BS oplev table) had been unplugged, which was unexpected.  I finally figured out that that's what the problem was with half of the accelerometers, plugged them back in, and now all of the sensors are up and running.

TheSEIS_GUR seismometer is under MC1, and all the others (the other Guralp, the Ranger which is oriented vertically, and all 6 accelerometers) are under MC2.

ELOG V3.1.3-